# Floquet-Lindblad Master Equation Approach to Open Quantum System Dynamics

### v1

@@ I Introduction
Computational physics has been a backbone of experimental and theoretical physics since the 1940’s, when the first nuclear bomb and ballistic simulations were performed at Los Alamos Laboratory [1]. Later on in the same decade, the first Monte Carlo methods were created, becoming one of the most ubiquitous and important simulation methods that is still in use today [2]. Later, in 1976, Goran Lindblad created what is now known as the Lindblad equation, opening a whole field of study into dissipative or “open” quantum systems [3]. Nowadays, simulation of open quantum systems has been formalized into coding packages and user-friendly functions, such as QuTiP or Qiskit, which allow users to simulate these systems without having the in-depth knowledge required to build the computational backbone from the ground-up [4, 5]. QuTiP specifically is described as an “open-source software package for the simulation of open quantum systems.” QuTiP provides several solvers with different regimes of validity, such as its MESolve function, which is a good first-pass simulation tool that utilizes the Lindblad equation to simulate open quantum systems. Also within QuTiP is FMMESolve, a solver that is ostensibly meant to be well-suited to the Floquet regime of ODEs, i.e., those with time-periodic coefficients.
@@
Here we describe and demonstrate a new solver named FLiMESolve that is suited for the Floquet regime of ordinary differential equations. This solver avoids some of the pitfalls and complications of MESolve and FMMESolve. First, it is specifically designed for time-periodic Hamiltonians, which leads to a solver that is multiple times faster than MESolve for problems in the Floquet regime. Furthermore, FLiMESolve makes use of the more empirically motivated decay rates, as opposed to spectral density functions, making it more accessible than FMMESolve for beginners. Finally, as shown by simulations of example systems, FLiMESolve can relax the secular approximation, allowing it to deal with periodic Hamiltonians with a time dependence more complex than a single frequency while still keeping a high computational efficiency.
@@ II.1 Lindblad Master Equation with Floquet Operators
In the above and in the following work, $\hbar$ is taken to be one. Then, the Borne-Markov master equation describes the time evolution of the density matrix operator of the open system by tracing over the states of a bath:
@@
In the above, $\hat{H}_{I},\hat{\rho}_{S},\hat{\rho}_{B}$ refer to the interaction picture Hamiltonian and system and bath density matrices, respectively.
The first step in moving from the Born-Markov master equation to the Floquet-Lindbladian is to fully expand the system-bath Interaction Hamiltonian, $\hat{H}_{I}$. The decomposed interaction Hamiltonian $\hat{H}_{I}(t)$ is given by
@@
The time-evolution operator, defined as $U(t,t_{0})=U_{s}(t,t_{0})\otimes U_{b}(t,t_{0})=U_{s}(t,t_{0})\otimes e^{i\hat{H}_{B}t}$ can be used to move the system and bath operators to the interaction picture as
@@
$\tilde{B}(t)=e^{i\hat{H}_{B}t}Be^{-i\hat{H}_{B}t}$ (5)
Using this decomposition in the Born-Markov master equation (Eqn. 2, above) gives:
$\begin{split}\frac{d}{dt}\hat{\rho}_{S}=-tr_{B}\{[S(t)\otimes&B(t),\\ \int_{t_{0}}^{t}[S(t^{\prime})\otimes B(t^{\prime}),\hat{\rho}_{S}(t)&\otimes\hat{\rho}_{B}]dt^{\prime}]\}\end{split}$ (6)
Next, it’s assumed that the system Hamiltonian is periodic in time with some fundamental frequency $\omega$. In a real open quantum system, this is usually accomplished through the application of time varying electric or magnetic field, e.g., laser light or microwave fields. Floquet’s theorem asserts that the general form of the solution of a linear differential equation with coefficients that are periodic takes the form [8]
@@
In Eqn. (9), $N$ is the dimensionality of the system, and $\epsilon_{\beta}$ refers to a Floquet quasienergy of the system. Using the above, the unitary time-evolution operator for the system Hamiltonian can be rewritten as
$U(t,t_{0})=e^{-i\epsilon_{\beta}t}\sum_{\beta}\ket{\phi_{\beta}(t)}\bra{\phi_{\beta}(0)}=\sum_{\beta}\ket{\Psi_{\beta}(t)}\bra{\Psi_{\beta}(0)}$ | (10)
@@
A single index pair ($\alpha,\beta$) in the equation above defines a single matrix element of the operator in the Floquet basis as
$S_{\alpha,\beta}(t)=$ $\bra{\Psi_{\alpha}(t)}\hat{S}\ket{\Psi_{\beta}(t)}$ (12)
To get to the final form of the operators used in the present work, two more operations are necessary. First, the periodicity of the field-interaction Hamiltonian means that the expectation value in Eqn. 12 can be re-expressed using a Fourier series of harmonics of the system frequency
@@
Secondly, operators are explicitly taken to be in the Floquet mode basis. This means that a new operator $S_{\alpha,\beta}^{k}(t)$ can be defined as
$S_{\alpha,\beta}^{k}(t)=e^{i\Delta_{\alpha,\beta}^{(k)}\omega t}\mathcal{S}_{\alpha\beta}(k)\ket{\phi_{\alpha}(0)}\bra{\phi_{\beta}(0)}$ | (14)
@@
Returning now to the Born-Markov master equation, the outer commutator of Eqn. 2 can be expanded to yield:
@@
$+$ $\left[\hat{S}(t),\hat{\rho}_{S}(t)\hat{S}^{\dagger}(t^{\prime})\right]Tr_{B}\left(\hat{B}^{\dagger}(t^{\prime})\hat{B}(t)\hat{\rho}_{B}\right)$
@@
With these operators, and some time-consuming algebra, the equation of motion can once again be written in a Lindblad form. Defining the system Hamiltonian as $\hat{H}_{S}$, the Lamb shift Hamiltonian as $\hat{H}_{LS}$,the system density matrix as $\rho_{S}$, and the dissipator as $D(\rho_{S})$, the Lindblad equation is given as [14]
$\hat{\dot{\rho}}=-i[\hat{H}_{S}+\hat{H}_{LS},\hat{\rho}_{S}]+D(\rho_{S})$ (17)
@@
$\hat{D}(\hat{\rho}_{S})=$ $\left[L^{1}(t)\rho_{S}(t)L^{1^{\dagger}}(t)-\frac{1}{2}\{L^{1\dagger}(t)L^{1}(t),\rho_{S}(t)\}\right]$ (18)
-- $\left[L^{0}(t)\rho_{S}(t)L^{0^{\dagger}}(t)-\frac{1}{2}\{L^{0\dagger}(t)L^{0}(t),\rho_{S}(t)\}\right]$
-- $\left[S(t)\rho_{S}(t)S^{\dagger}(t)-\frac{1}{2}\{S^{\dagger}(t)S(t),\rho_{S}(t)\}\right]$
$\hat{H}_{LS}=$ $\frac{1}{2}[\hat{L}^{0^{\dagger}}(t)\hat{S}(t)-\hat{S}^{\dagger}(t)L^{0}(t)]$
Using the previously defined Floquet decomposition on the system operators, the operators in Eqn. 16 can be re-expressed as
$L^{0,k}_{\alpha,\beta}(t)=\sum_{\alpha\beta}\Gamma_{\alpha,\beta}^{k}(t)S_{\alpha,\beta}^{k}(t),\quad$ (19)
$L^{1,k}_{\alpha,\beta}(t)=\sum_{\alpha\beta}\left(\Gamma_{\alpha,\beta}^{k}(t)+1\right)S_{\alpha,\beta}^{k}(t),$
@@
Finally, a Markov approximation is made for the bath autocorrelation functions. The autocorrelation function in the last line of Eqn. 19 can be understood as a Fourier transform. By assuming that the bath correlation times are effectively instant (i.e. making the assumption that the spectral density function of the bath takes the form of white noise), the decay rate of the Lindblad terms in the master equation becomes constant:
@@
$\begin{gathered}\hat{D}=\sum_{\alpha,\alpha^{\prime}}\sum_{\beta,\beta^{\prime}}\gamma[S_{\alpha,\beta}^{k}(t)\rho(t)S_{\alpha^{\prime},\beta^{\prime}}^{k^{\prime\dagger}}(t))\\ -\frac{1}{2}S_{\alpha^{\prime},\beta^{\prime}}^{k^{\prime\dagger}}(t)S_{\alpha,\beta}^{k}(t)\rho(t)-\frac{1}{2}\rho(t)S_{\alpha^{\prime},\beta^{\prime}}^{k^{\prime\dagger}}(t)S_{\alpha,\beta}^{k}(t)]\end{gathered}$ (21)
II.2 The Floquet-Lindblad Rate Matrix R(t)
@@
The final step in this process of finding a computationally-efficient form the Floquet-Lindblad master equation is to solve for the individual matrix elements of equation (21) and rearranging them into a rate matrix form. A convenient basis to use for this operation is the Floquet mode basis at time $t=0$, as discussed previously and which led to Eqn. (14). Making this transformation, and after another round of time-consuming algebra, the the form of the master equation given by $\frac{d}{dt}\hat{\rho}_{S}(t)=\hat{R}(t)\hat{\rho}_{S}(t)$ is obtained, with the matrix $\hat{R}(t)$ having elements given by:
@@
$R{}_{mn,pq}(t)=\sum_{\alpha\beta\alpha^{\prime}\beta^{\prime}}\sum_{kk^{\prime}=-\infty}^{\infty}$ $S_{\alpha\beta}(k)S_{\alpha^{\prime}\beta^{\prime}}^{*}\left(k^{\prime}\right)e^{i\left(\Delta_{\alpha\beta}^{(k)}-\Delta_{\alpha^{\prime}\beta^{\prime}}^{(k^{\prime})}\right)\omega_{0}t}$ (22)
@@
$*\left[\delta_{n\alpha}\delta_{\alpha^{\prime}m}\delta_{p\beta}\delta_{q\beta^{\prime}}-\frac{1}{2}\left(\delta_{n\beta^{\prime}}\delta_{\alpha^{\prime}\alpha}\delta_{p\beta}\delta_{qm}+\delta_{m\beta^{\prime}}\delta_{\alpha^{\prime}\alpha}\delta_{pn}\delta_{q\beta^{\prime}}\right)\right]$
@@
Note that in equation (22), the rate matrix $R_{mn,pq}(t)$ is four-indexed, but only two dimensional. In order to obtain the first-order ODE form of the master equation, it is necessary to “unfold” the density matrices into one-dimensional supervectors. Given that the density matrix $\hat{\rho}$ is indexed by $n,m$, and that the dimensionality of the system is (again) given by $N$, the unfolding process is given as
$\begin{gathered}\hat{\rho}_{n,m}=\vec{\rho}_{n+N*m}\end{gathered}$ (23)
@@
In Eqn. 23, $\hat{\rho}$ explicitly refers to the matrix form of the density matrix, and $\vec{\rho}$ refers to the supervector form of the density matrix. Using this notation, the first-order ODE given by equation (22) can be written with Einstein notation as
$\begin{gathered}\dot{\vec{\rho}}_{(n+B\times m)}=\hat{R}_{(m+B\times n),(p+B\times q)}(t)\vec{\rho}_{(p+B\times q)}\end{gathered}$ (24)
@@
The above form of the Master Equation is then solvable as a first-order ODE using QuTiP’s ODE infrastructure to evolve suitable systems forward in time in an efficient manner. This efficiency lends itself well to systems with long-time dissipation dynamics but short-time system dynamics, and an example is shown below. Additionally, the time-dependent nature of equation (24) allows for the relaxation of the secular approximation, which will be covered in the other example below.

### v2

@@ I Introduction
Computational physics has been a backbone of experimental and theoretical physics since the 1940’s, when the first nuclear bomb and ballistic simulations were performed at Los Alamos Laboratory [1]. Later on in the same decade, the first Monte Carlo methods were created, becoming one of the most ubiquitous and important simulation methods that is still in use today [2]. Later, in 1976, Goran Lindblad created what is now known as the Lindblad equation, opening a whole field of study into dissipative or “open” quantum systems [3]. Nowadays, simulation of open quantum systems has been formalized into coding packages and user-friendly functions, such as QuTiP or Qiskit, which allow users to simulate these systems without having the in-depth knowledge required to build the computational backbone from the ground-up [4, 5, 6]. QuTiP specifically is described as an “open-source software package for the simulation of open quantum systems.” QuTiP provides several solvers with different regimes of validity, such as its MESolve function, which is a good first-pass simulation tool that utilizes the Lindblad equation to simulate open quantum systems. Also within QuTiP is FMMESolve, a solver that is ostensibly meant to be well-suited to the Floquet regime of ODEs, i.e., those with time-periodic coefficients.
@@
Here we describe and demonstrate a new solver for QuTiP named FLiMESolve that is suited for the Floquet regime of ordinary differential equations. This solver avoids some of the pitfalls and complications of MESolve and FMMESolve. First, it is specifically designed for time-periodic Hamiltonians, which leads to a solver that is multiple times faster than MESolve for problems in the Floquet regime. Furthermore, FLiMESolve makes use of the more empirically motivated decay rates, as opposed to spectral density functions, making it more accessible than FMMESolve for beginners. Finally, as shown by simulations of example systems, FLiMESolve can go beyond the secular approximation, allowing it to deal with periodic Hamiltonians with a time dependence more complex than a single frequency while still keeping a high computational efficiency.
@@ II.1 Lindblad Master Equation with Floquet Operators
For brevity, in the following, $\hbar$ is taken to be one. The Born-Markov master equation describes the time evolution of the density matrix operator of the open system by tracing over the states of the bath:
@@
In the above, $\hat{H}_{I}$, $\hat{\rho}_{S}$, $\hat{\rho}_{B}$ refer to the interaction picture Hamiltonian and system and bath density matrices in the interaction picture, respectively. Further, the time $t_{0}$ is some initial time of the system. Additionally, the interaction picture itself is interpreted to be the Floquet basis, i.e. moving to the interaction picture is the same as moving to the Floquet basis.
The first step in moving from the Born-Markov master equation to the Floquet-Lindblad master equation is to fully expand the system-bath interaction Hamiltonian, $\hat{H}_{I}$. The decomposed interaction Hamiltonian in the interaction picture $\hat{H}_{I}(t)$ is given by the tensor product of the system operators, $\hat{S}(t)$, and the bath operators, $\hat{B}(t)$, as
@@
The unitary time-evolution operator which evolves the system from time $t_{0}$ to time $t$ and without dissipation, defined as $U(t,t_{0})=U_{s}(t,t_{0})\otimes U_{b}(t,t_{0})=U_{s}(t,t_{0})\otimes e^{i\hat{H}_{B}t}$ can be used to move the system operators $\hat{S}$ and bath operators $\hat{B}$ to the interaction picture as
@@
$\hat{B}(t)=e^{i\hat{H}_{B}t}\hat{B}e^{-i\hat{H}_{B}t}$ (5)
Using this decomposition in the Born-Markov master equation (Eqn. (2), above) gives:
$\begin{split}\frac{d}{dt}\hat{\rho}_{S}=-\operatorname{Tr}_{B}\{[S(t)\otimes&B(t),\\ \int_{t_{0}}^{t}[S(t^{\prime})\otimes B(t^{\prime}),\hat{\rho}_{S}(t)&\otimes\hat{\rho}_{B}]dt^{\prime}]\}\end{split}$ (6)
Next, it’s assumed that the system Hamiltonian is periodic in time with some fundamental frequency $\omega$. In a real open quantum system, this is usually accomplished through the application of time varying electric or magnetic field, e.g., laser light or microwave fields. Floquet’s theorem asserts that the general form of the solution of a linear differential equation, like the Schrödinger equation, with coefficients that are periodic takes the form [7]
@@
In Eqn. (9), $N$ is the dimensionality of the system, and $\epsilon_{\beta}$ refers to a Floquet quasienergy of the system. These quasinergies are the entries of the diagonal matrix defined in Eq. (7). Using the above, the unitary time-evolution operator for the system Hamiltonian can be rewritten as
$U(t,t_{0})=\sum_{\beta}\ket{\Psi_{\beta}(t)}\bra{\Psi_{\beta}(0)}=e^{-i\epsilon_{\beta}t}\sum_{\beta}\ket{\phi_{\beta}(t)}\bra{\phi_{\beta}(0)}$ | (10)
@@
To get to the final form of the operators used in the present work, two more operations are necessary. First, the periodicity of the field-interaction Hamiltonian means that the expectation value in Eqn. (11) can be re-expressed using a Fourier series of harmonics of the system frequency
@@
In the above equation, the coefficient $\mathcal{S}_{\alpha,\beta}(k)$ is the $k^{th}$ Fourier coefficient of the $\alpha,\beta$ matrix element of the system operator in the interaction picture. Secondly, operators are explicitly taken to be in the Floquet mode basis. This means that a new operator $\hat{S}_{\alpha,\beta}^{k}(t)$ can be defined as
$\hat{S}_{\alpha,\beta}^{k}(t)=e^{i\Delta_{\alpha,\beta}^{(k)}\omega t}\mathcal{S}_{\alpha\beta}(k)\ket{\phi_{\alpha}(0)}\bra{\phi_{\beta}(0)}$ | (13)
@@
Returning now to the Born-Markov master equation, the outer commutator of Eqn. (2) can be expanded to yield:
@@
$+$ $\left[\hat{S}(t),\hat{\rho}_{S}(t)\hat{S}^{\dagger}(t^{\prime})\right]\operatorname{Tr}_{B}\left\{\hat{B}^{\dagger}(t^{\prime})\hat{B}(t)\hat{\rho}_{B}\right\}$
@@
With these operators, and some time-consuming algebra, the equation of motion (14) can once again be written in a Lindblad form. Defining the system Hamiltonian as $\hat{H}_{S}$, the Lamb shift Hamiltonian as $\hat{H}_{LS}$,the system density matrix as $\hat{\rho}_{S}$, and the dissipator as $\hat{D}(\hat{\rho}_{S})$, the Lindblad equation is given as [15]
$\frac{d}{dt}\hat{\rho}=-i[\hat{H}_{S}+\hat{H}_{LS},\hat{\rho}_{S}]+\hat{D}(\hat{\rho}_{S})$ (16)
@@
$\hat{D}(\hat{\rho}_{S})=$ $\left[\hat{L}^{1}(t)\hat{\hat{\rho}}_{S}(t)\hat{L}^{1\dagger}(t)-\frac{1}{2}\{\hat{L}^{1\dagger}(t)\hat{L}^{1}(t),\hat{\rho}_{S}(t)\}\right]$ (17)
-- $\left[\hat{L}^{0}(t)\hat{\rho}_{S}(t)\hat{L}^{0^{\dagger}}(t)-\frac{1}{2}\{\hat{L}^{0\dagger}(t)\hat{L}^{0}(t),\hat{\rho}_{S}(t)\}\right]$
-- $\left[\hat{S}(t)\hat{\rho}_{S}(t)\hat{S}^{\dagger}(t)-\frac{1}{2}\{\hat{S}^{\dagger}(t)\hat{S}(t),\hat{\rho}_{S}(t)\}\right]$
$\hat{H}_{LS}=$ $\frac{1}{2}[\hat{L}^{0{\dagger}}(t)\hat{S}(t)-\hat{S}^{\dagger}(t)\hat{L}^{0}(t)]$
Using the previously defined Floquet decomposition on the system operators, the operators in Eqn. (15) can be re-expressed as
$\hat{L}^{0,k}_{\alpha,\beta}(t)=\sum_{\alpha\beta}\Gamma_{\alpha,\beta}^{k}(t)\hat{S}_{\alpha,\beta}^{k}(t),\quad$ (18)
$\hat{L}^{1,k}_{\alpha,\beta}(t)=\sum_{\alpha\beta}\left(\Gamma_{\alpha,\beta}^{k}(t)+1\right)\hat{S}_{\alpha,\beta}^{k}(t),$
@@
Finally, a Markov approximation is made for the bath autocorrelation functions. The autocorrelation function in the last line of Eqn. 18 can be understood as a Fourier transform. By assuming that the bath correlation times are effectively instant (i.e. making the assumption that the spectral density function of the bath takes the form of white noise), the decay rate of the Lindblad terms in the master equation becomes constant:
@@
$\begin{gathered}\hat{D}(\hat{\rho}_{S})=\sum_{\alpha,\alpha^{\prime}}\sum_{\beta,\beta^{\prime}}\gamma[\hat{S}_{\alpha,\beta}^{k}(t)\hat{\rho}(t)\hat{S}_{\alpha^{\prime},\beta^{\prime}}^{k^{\prime\dagger}}(t))\\ -\frac{1}{2}\hat{S}_{\alpha^{\prime},\beta^{\prime}}^{k^{\prime\dagger}}(t)\hat{S}_{\alpha,\beta}^{k}(t)\hat{\rho}(t)-\frac{1}{2}\hat{\rho}(t)\hat{S}_{\alpha^{\prime},\beta^{\prime}}^{k^{\prime\dagger}}(t)\hat{S}_{\alpha,\beta}^{k}(t)]\end{gathered}$ (20)
II.2 The Floquet-Lindblad Rate Matrix, R(t)
@@
The final step in the process of finding a computationally-efficient form of the Floquet-Lindblad master equation is to find an expression for the individual matrix elements of Eqn. (20) and rearranging them into a rate matrix form. In order to obtain the first-order ODE form of the master equation, it is necessary to “unfold” the density matrices into one-dimensional supervectors. A convenient basis to use for this operation is the Floquet mode basis at time $t=0$, as discussed previously and which led to Eqn. (13). Making this transformation, and after another round of time-consuming algebra, the the form of the master equation given by $\frac{d}{dt}\hat{\rho}_{S}(t)=\hat{R}(t)\hat{\rho}_{S}(t)$ is obtained, with the matrix $\hat{R}(t)$ having elements given by:
@@
$R{}_{mn,pq}(t)=\sum_{\alpha\beta\alpha^{\prime}\beta^{\prime}}\sum_{kk^{\prime}=-\infty}^{\infty}$ $S_{\alpha\beta}(k)S_{\alpha^{\prime}\beta^{\prime}}^{*}\left(k^{\prime}\right)e^{i\left(\Delta_{\alpha\beta}^{(k)}-\Delta_{\alpha^{\prime}\beta^{\prime}}^{(k^{\prime})}\right)\omega t}$ (21)
@@
×$\times\left[\delta_{n\alpha}\delta_{\alpha^{\prime}m}\delta_{p\beta}\delta_{q\beta^{\prime}}-\frac{1}{2}\left(\delta_{n\beta^{\prime}}\delta_{\alpha^{\prime}\alpha}\delta_{p\beta}\delta_{qm}+\delta_{m\beta^{\prime}}\delta_{\alpha^{\prime}\alpha}\delta_{pn}\delta_{q\beta^{\prime}}\right)\right]$
@@
Note that in equation (21), the rate matrix $R_{mn,pq}(t)$ is four-indexed, but only two dimensional. Given that the density matrix $\hat{\rho}$ is indexed by $n,m$, and that the dimensionality of the system is (again) given by $N$, the unfolding process is given as
$\begin{gathered}\hat{\rho}_{n,m}=\vec{\rho}_{n+Nm}\end{gathered}$ (22)
@@
In Eqn. (22), $\hat{\rho}$ explicitly refers to the matrix form of the density matrix, and $\vec{\rho}$ refers to the supervector form of the density matrix. Using this notation, the first-order ODE given by equation (21) can be written with Einstein notation as
$\begin{gathered}\dot{\vec{\rho}}_{(n+Bm)}=\hat{R}_{(m+Bn),(p+Bq)}(t)\vec{\rho}_{(p+Bq)}\end{gathered}$ (23)
@@
The above form of the Master Equation is then solvable as a first-order ODE using QuTiP’s ODE infrastructure to evolve suitable systems forward in time in an efficient manner. This efficiency lends itself well to systems with long-time dissipation dynamics but short-time system dynamics. Additionally, the time-dependent nature of Eqn. (23) allows for the relaxation of the secular approximation. Several examples are shown in the next section to illustrate these advantages.
II.3 Secular Approximation and the Negligibility Factor
Eqn. 21 contains within it all time-dependent and independent dynamics of any given system. However, calculating the solution to the full time-dependent Floquet-Lindblad equation can be extremely time consuming and is unnecessary for many applications. For the sake of efficiency, the so-called ”negligibility factor” is used. This negligibility factor is defined as:
$\mathbb{N}_{\alpha,\alpha^{\prime},\beta,\beta^{\prime}}^{k,k^{\prime}}=\frac{\left(\Delta_{\alpha\beta}^{k}-\Delta_{\alpha^{\prime}\beta^{\prime}}^{k^{\prime}}\right)\omega}{\mathcal{S}_{\alpha\beta}^{(k)}\mathcal{S}_{\alpha^{\prime}\beta^{\prime}}^{(k^{\prime})}}$ (24)
The numerator of Eq. (24) is the frequency difference found in the exponential term of Eq. (21). This term, with $\Delta_{\alpha\beta}^{k}=\left(\frac{\epsilon_{\alpha}}{\omega}-\frac{\epsilon_{\beta}}{\omega}+k\right)$ as the difference in rotation frequency of the two three-indexed system operator terms. This difference in frequencies is itself a frequency, which defines how quickly a term in Eq. (21) rotates through the complex plane. If a term rotates too quickly, then the effect of the rate product is effectively averaged out over any significant period of time.
The denominator of Eq. (21), which is the product of two three-indexed system operators decomposed along a single frequency (with the harmonic multiple denoted by $k$), and two state indices ($\alpha$ and $\beta$ denote the Floquet modes by which the operator is sandwiched), describes the magnitude of the effect of the full Fourier-expanded collapse operator product (which is indexed by $\alpha,\alpha^{\prime},\beta,\beta^{\prime},k,k^{\prime}$) on the system. A larger magnitude has a larger affect on the system, and a smaller magnitude the converse.
The effect of the quotient is to determine which terms of the overall Floquet-Lindblad master equation are ”important” enough to be kept. A low value for the negligibility factor means that either the frequency is low or the rate product is large, corresponding to a term that appreciably affects the Hamiltonian on appreciable timescales. Conversely, a large value in the negligibility factor means that either the term rotates very quickly or that the rate product is small, leading to a term that will average out on appreciable timescales. An example of where the negligibility factor becomes important can be seen in the final example system below.